OpenStudy (xcecex):

Please help! I only have 20 min :( I would appreciate it one of you kind fellows would help me:) Also a medal and fan will be awarded :) My question -4.5[2x-(0.5-1.4)] A.-9x+0.9 B.-9x-0.9 C-9x-4.05 D9x-4.05

OpenStudy (xcecex):

Simplify^.

OpenStudy (anonymous):

Just pemdas it out

OpenStudy (anonymous):

Do the things in the bracket first.

OpenStudy (xcecex):

Okaaay so,0.5-1.4 first?

OpenStudy (anonymous):

-4.5[2x-(-0.9)]

OpenStudy (anonymous):

-4.5[2x+ 0.9]

OpenStudy (anonymous):

-9x - 4.05

OpenStudy (anonymous):

.5-1.4=-0.9, so the expression becomes -4.5(2x-(-.9))=-4.5(2x+0.9) Then just expand the equation and you'll get the answer (so do -4.5*2x+-4.5*0.9)

OpenStudy (xcecex):

Okay hold me lemmie calculate dat >_<

OpenStudy (xcecex):

okay so i got -9x+-4.05

OpenStudy (xcecex):

@Fzxt

OpenStudy (xcecex):

@swagmaster47

OpenStudy (anonymous):

Yep. This is the same as -9x - 4.05, which is option C.

OpenStudy (xcecex):

I thought it would be harder dan that O.o

OpenStudy (xcecex):

Okay so i have 3 more questions :/